# Expert Consensus-based Video-Based Assessment Tool for Workflow Analysis in Minimally Invasive Colorectal Surgery: Development and Validation of ColoWorkflow

### v1

@@ 3.1 Modified Delphi study for expert consensus on colorectal workflow descriptors
Twenty experts attended the final consensus meeting. Remaining items were reviewed, discussed, and revised for clarity and precision. Key modifications included splitting the phase “Other Interventions” into 2 distinct phases: “Preplanned Additional Procedures” and “Unplanned Procedures”. The definition of “Leak Testing” was modified, and the description of “Mesorectal Dissection” was refined to specify circumferential dissection with or without vascular control. Following these revisions, expert consensus was achieved on 10 procedural phases and 34 workflow steps (Figure 2) and their description. Complete phase and step definitions are detailed in Supplementary Tables 3 and 4.
@@ 6.2 Funding Statement
1 Insertion of ports and instruments (intraabdominal part) Trocar entry into the abdominal cavity and eventual insertion of instruments.
@@
2 Abdominal cavity and structures assessment Visual/instrumental assessment of bowel, liver, peritoneum, and other structures.
@@
3 Adhesiolysis/structure division for exposure Dividing adhesions and/or falciform ligament to delineate anatomy and optimize exposure of the target area.
@@
4 Mesentery/mesocolon exposure Lifting and repositioning omentum, small bowel, or colon to expose the mesentery/mesocolon root.
@@
5 Peritoneum incision and Toldt’s/Gerota fascia dissection – left Incision of the peritoneum with subsequent dissection to isolate key vessels and develop the retroperitoneal plane between the Toldt and Gerota fasciae.
@@
Fig. 2: Inter-annotator differences between the annotators and distribution across the labels. The plots show the Cohen’s $\kappa$ metric and its variability for each label, indicating that some labels have high agreement with low variation (e.g., step 1), while others may have low agreement and high variability (e.g., step 20). Phases shown in (a) (above); Steps in (b) (below). Phase and step names correspond to the numbering given in eTables 3 and 4.

### v2

@@ 3.1 Modified Delphi study for expert consensus on colorectal workflow descriptors
Twenty experts attended the final consensus meeting. Remaining items were reviewed, discussed, and revised for clarity and precision. Key modifications included splitting the phase “Other Interventions” into 2 distinct phases: “Preplanned Additional Procedures” and “Unplanned Procedures”. The definition of “Leak Testing” was modified, and the description of “Mesorectal Dissection” was refined to specify circumferential dissection with or without vascular control. Following these revisions, expert consensus was achieved on 10 procedural phases and 34 workflow steps (Figure 2) and their description. Complete phase and step definitions are detailed in Supplementary Tables 3 and LABEL:etab:4.
@@ 6.2 Funding Statement
1 Insertion of ports and instruments (intraabdominal part) Trocar entry into the abdominal cavity and eventual insertion of laparoscopic instruments.
@@
2 Abdominal cavity and structures assessment Visual and instrumental assessment of bowel, liver, peritoneum, and other intraabdominal structures.
@@
3 Adhesiolysis/structure division for exposure Division of adhesions and/or falciform ligament to delineate anatomy and optimize exposure.
@@
4 Mesentery/mesocolon exposure Lifting and repositioning omentum, small bowel, or colon to expose the mesentery or mesocolon root.
@@
5 Peritoneum incision and Toldt’s/Gerota fascia dissection – left Peritoneal incision with dissection to isolate inferior mesenteric vessels and develop the retroperitoneal plane between Toldt and Gerota fasciae.
6 Peritoneum incision and Toldt’s/Gerota/Fredet fascia dissection – right Peritoneal incision with retroperitoneal dissection to isolate relevant vessels, visualize the duodenum, and develop planes between Toldt, Gerota, and Fredet fasciae.
7 Inferior mesenteric artery isolation, ligation and division – left Dissection, ligation, and division of the inferior mesenteric artery.
8 Inferior mesenteric vein isolation, ligation and division – left Identification, dissection, ligation, and division of the inferior mesenteric vein near the ligament of Treitz.
9 Ileocolic vessels isolation, ligation and division – right Dissection, ligation, and division of the ileocolic artery and vein.
10 Middle colic and/or Henle trunk branches isolation, ligation and division – right Dissection, ligation, and division of middle colic branches with or without right colic artery and Henle trunk branches.
11 Mesosigmoid/mesocolon/mesentery division Division of mesosigmoid, mesocolon, or mesentery parallel to bowel wall or along vascular pedicles in preparation for bowel transection.
12 Sigmoid mobilisation and lateral dissection – left Lateral mobilization of sigmoid and descending colon by dividing the lateral peritoneal reflection along the Monk line.
13 Caecal mobilisation and lateral dissection – right Lateral mobilization of caecum and ascending colon by dividing inferior and lateral peritoneal attachments along the Monk line.
14 Lesser sac entry and/or omentum division Division of gastrocolic ligament or colo-epiploic attachments to enter the lesser sac and/or divide omentum.
15 Splenic flexure mobilisation Division of splenocolic, pancreaticocolic, and phrenocolic ligaments with mobilization of the splenic flexure.
16 Hepatic flexure mobilisation Division of hepatic flexure attachments to mobilize the right colonic flexure.
17 Mesorectum dissection Circumferential dissection of the mesorectum with or without vascular control down to the planned transection level.
18 Distal resection site selection and preparation – left Circumferential clearing of pericolonic or perirectal fat at the planned distal resection site with optional perfusion assessment.
19 Rectum/sigmoid transection – left Transection of the distal bowel segment at the planned resection level using stapling devices.
20 Proximal resection site preparation and transection (intraabdominal part) Intraabdominal preparation and marking of the proximal bowel with circumferential clearing for planned transection.
21 Transected bowel handling and externalization (intraabdominal part) Intraabdominal preparation for bowel externalization including incision, wound protector placement, and manipulation.
22 Distal resection site preparation and transection – right Preparation and transection of the distal bowel end for right-sided resections with optional perfusion assessment.
23 Ileal preparation and transection Preparation and transection of the proximal ileal segment for right-sided resections with optional perfusion assessment.
24 Dye injection and visualization Injection of dye to assess bowel perfusion and guide resection margins.
25 Ileoanal pouch creation (intraabdominal part) Construction of an ileal pouch for ileoanal anastomosis with attachment of the stapler anvil.
26 Preparation for anastomosis per rectum – left Advancement and positioning of the proximal bowel into the pelvis for tension-free colorectal anastomosis.
27 Rectal stump perforation and stapler firing Perforation of rectal stump with stapler spike followed by stapler approximation and firing.
28 Leak testing – left Assessment of anastomotic integrity using air, saline, dye, and/or endoscopic evaluation.
29 Preparation for intracorporeal anastomosis – right Alignment and positioning of proximal and distal bowel ends for intracorporeal anastomosis.
30 Enterotomy, colotomy and intracorporeal anastomosis – right Creation of enterotomy and colotomy followed by stapled intracorporeal anastomosis and bleeding assessment.
31 Enterotomy/colotomy closure – right Closure of the common enterotomy to restore bowel continuity.
32 Additional suturing, stapling or suture removal Reinforcement of anastomosis, closure of mesenteric defects, and/or removal of temporary sutures.
33 Stoma creation (intraabdominal part) Preparation and positioning of bowel for stoma creation with verification of orientation and tension.
34 Washing, aspiration, drain insertion, and port removal (intraabdominal part) Irrigation, aspiration, hemostasis, drain placement if required, and removal of trocars and ports.
Table 4: Video-based assessment tool: list of steps and their descriptions reaching consensus after all rounds of Delphi. (continued)
@@
Fig. 2: Inter-annotator differences between the annotators and distribution across the labels. The plots show the Cohen’s $\kappa$ metric and its variability for each label, indicating that some labels have high agreement with low variation (e.g., step 1), while others may have low agreement and high variability (e.g., step 20). Phases shown in (a) (above); Steps in (b) (below). Phase and step names correspond to the numbering given in eTables 3 and LABEL:etab:4.